Answers the questions
34 OHAPTER COST AND SALES CONCEPTS c. The banquet manager is planning a function for 120 persons for next Friday evening Each guest will be given one glass of wine. How many bottles should be ordered for the party? d. What will be the unit cost of the wine? The total cost? 10 Sales records for a luncheon in the Newmarket Restaurant for a recent week were Item A, 196 Item B, 72 Item C, 142 Item D, 24 ltem E, 112 Item F.224 Item G, 162 Given this information, calculate the sales mix. 11.) Calculate the average check from the following data: a. Sales, $1.000.00; Number of customers, 125 b. Sales, $1.300.00: Number of customers, 158 c. Sales, $8,720.53 Number of customers, 976 The following table indicates the number of covers served and the gross sales per server for one three-hour period in Sallys Restaurant. Determine: (a) the average number of covers served per hour per server, and (b) the average sale per server for the three-hour period. Covers Served 71 Gross Sales Per Server $237.40 $263.95 S188.25 Server 58 3) Use the information about Sallys Restaurant identfied in Question 12 i Use the informnation about Sals Restaurant identifed Question12 to complete the following a. Calculate the average check b. Calculate the turnover for the three-hour period if there are 65 seats in the restaurant. (14. Given the information about Sallys Restaurant identified in Questions 12 and 13, assume the restaurant had 85,629 customers per year an gross sales were $352,783.40 a. Calculate the average check b. Calculate sales per seat for the year
Answer #1

Problem 10:
In order to compute the sales mix, one needs to compute the total sales, and then compute the percentage sales for each product.
Total sales = 196 + 72 + 142 + 24 + 112 + 224 + 162 = 932

Then, for each item, we compute its percentage sales with respect to total.

Item A: 196/932 * 100 = 21.03%
Item B: 72/932 * 100 = 7.73%
Item C: 142/932 * 100 = 15.24%
Item D: 24/932 * 100 = 2.58%
Item E: 112/932 * 100 = 12.02%
Item F: 224/932 * 100 = 24.03%
Item G: 162/932 * 100 = 17.38%

Problem 11:
To compute average check, you simply need to divide sales by the number of customers.
a. $1,000/125 = $8
b. $1,300/158 = $8.23
c. $8720.52/976 = $8.93

Problem 12:
a) To compute the average number of covers per server, one needs to compute the total number of covers, divided by the number of servers-hours.
Total cover = 71 + 66 + 58 = 195
Total server-hours = 9
Number of covers per server per hour = 195/9 = 21.33
b) Average sales per server = total sales / number of servers
= ($237.40 + $263.95 + $188.25)/3
= $229.87

Problem 13:
a. This is total gross sales / total number of covers
= ($237.40 + $263.95 + $188.25)/(71 + 66 + 58)
= $3.54

b. The word turnover is ambiguous. One meaning of turnover is the number of customers per seat. This would be
Number of customers (covers) / number of seats
= (71 + 66 + 58)/65
= 3 customers per seat.


Problem 14:
a) Average check = Total sales/total covers
= $352783.40/85629
= $4.12
b) Sales per seat = Total sales / number of seats
= $352783.40/65 = $5427.44 per seat.